OpenStudy (anonymous):

Change to a simplified fraction or mixed number. 0.4%

OpenStudy (anonymous):

2 over 5 is the answer correct?

OpenStudy (unklerhaukus):

0.40%==40/100=4/10=2/5

OpenStudy (anonymous):

4/10 = 2/5